OpenStudy (aaronandyson):

Evaluate log20 + log50 - 2log5 - 2log2

OpenStudy (aaronandyson):

@Kainui

OpenStudy (kainui):

Show me what you can do so I can help you where you're having trouble.

OpenStudy (aaronandyson):

log 20 x 50 x 4/25

OpenStudy (aaronandyson):

i had got log 20 x 50/25/4

OpenStudy (aaronandyson):

So as the final answer I got log160

OpenStudy (aaronandyson):

@Kainui

OpenStudy (kainui):

That looks pretty good, except you should have: \[\log \left( \frac{20*50}{4*25} \right)\] since all the ones that are subtracted should be divided and the ones that are added are multiplied on top.

OpenStudy (aaronandyson):

Log 25?

OpenStudy (aaronandyson):

sorry so now its log 5 x 2 so log 10 which is 1

OpenStudy (kainui):

Yeah, perfect!